# 2d Conformal Field Theories on Magic Triangle

### v1

@@ Abstract
The magic triangle due to Cvitanović and Deligne–Gross is an extension of the Freudenthal–Tits magic square of semisimple Lie algebras. In this paper, we identify all 2d rational conformal field theories associated to the magic triangle. These include various Wess–Zumino–Witten (WZW) models, Virasoro minimal models, compact bosons and their non-diagonal modular invariants. At level one, we find a two-parameter family of modular linear differential equation of fourth order whose solutions produce the affine characters of all elements in the magic triangle. We find a universal coset relation for the whole triangle which generalizes the dual pairs with respect to $(E_{8})_{1}$ in the Cvitanović–Deligne exceptional series. This leads to the dimension and degeneracy of each primary field and also to five atomic models which constitute all theories in the triangle. At level two, we find a special row of the triangle – the subexceptional series has novel $N=1$ supersymmetry, and the Neveu–Schwarz/Ramond characters satisfy a one-parameter family of fermionic modular linear differential equations. Moreover, we find many new coset constructions involving WZW models at higher levels.
@@ 1 Introduction
The magic triangle of groups was discovered in both theoretical physics Cvitanovic (1981) and mathematics Deligne and Gross (2002). It may be viewed as a triangular extension of the Freudenthal–Tits magic square $\mathfrak{m}(\mathbb{A}_{1},\mathbb{A}_{2})$ of semisimple Lie algebras and Lie groups, where $\mathbb{A}$ ranges over the four normed division algebras: the real numbers, complex numbers, quaternions, and octonions. From another perspective, the magic triangle can be regarded as a two-dimensional generalization of the celebrated Cvitanović–Deligne exceptional series Cvitanovic (1981, 2008); Deligne (1996). The original magic triangle of Lie groups introduced by Cvitanović (Cvitanovic, 1981, Table 6) is reproduced in Table 1. The magic triangle of Deligne and Gross appears in (Deligne and Gross, 2002, Table 1).
The magic triangle is symmetric, and each row (or equivalently, each column) is referred to as an exceptional series. In Table 1, the bottom row, terminating at $E_{8}$, is known as the Cvitanović–Deligne exceptional series. The row immediately above it, ending at $E_{7}$, is called the subexceptional series. The third row from the bottom, ending at $E_{6}$, is known as the Severi exceptional series, while the fourth row from the bottom, terminating at $F_{4}$, is referred to as the Severi-section exceptional series.
@@
Each exceptional series exhibits distinctive and remarkable properties in representation theory. For instance, the Cvitanović–Deligne exceptional series displays striking uniformity in the tensor product decompositions of adjoint representations, and the dimensions of its irreducible representations admit expressions as rational functions of the dual Coxeter number Deligne (1996); Cohen and de Man (1996); Deligne and de Man (1996). From the perspective of Vogel’s theory of the universal Lie algebra, both the Cvitanović–Deligne series and the subexceptional series correspond to straight lines in Vogel’s projective plane Vogel (1999). Moreover, the four exceptional series are deeply connected with geometry Landsberg and Manivel (2002, 2004, 2001). In particular, the terminology Severi exceptional series and Severi-section exceptional series reflects their geometric relationship with Severi varieties Landsberg and Manivel (2004).
@@
Several conceptual issues arise in the original magic triangle displayed in Table 1. First, despite its name, the original magic triangle of Lie groups does not form a complete triangular array, as certain entries along the slope are missing. These absent elements were identified as discrete groups in Deligne and Gross (2002). A second issue concerns the common practice in the physics literature of adjoining an additional entry “Lee–Yang model” at the beginning of the Cvitanović–Deligne exceptional series, which further obscures the triangular structure. A third question is whether the famous intermediate Lie algebra $E_{7+1/2}$ Landsberg and Manivel (2006), lying between $E_{7}$ and $E_{8}$, can be incorporated into the Cvitanović–Deligne exceptional series in a manner compatible with the magic triangle. In this work, we address these three issues in a unified framework and show that their resolution naturally requires an extension of the original diagram, resulting in a genuinely triangular structure.
@@
Our approach is based on two-dimensional rational conformal field theories (RCFTs) associated with the magic triangle. This may be viewed as a natural generalization of the work of Mathur, Mukhi, and Sen on the 2d RCFTs associated with the Cvitanović–Deligne exceptional series Mathur et al. (1988, 1989). From this perspective, the resulting structure can be interpreted as an affine analogue of the original magic triangle. The construction is particularly natural at level one; the resulting extended triangle is displayed in Table 2. For each Lie group appearing in the original magic triangle in Table 1, the corresponding entry in Table 2 is given by the associated Wess–Zumino–Witten model mostly at level one. The remaining entries consist of various minimal models, compact boson theories, and their non-diagonal modular invariants, which will be explained in detail below. In particular, we introduce an additional row ending at $E_{7+1/2}$, referred to as the intermediate exceptional series, which has been studied recently in Lee et al. (2024b).
@@ 0 $D_{2\rm A}$ $M_{2}$ $(A_{1}^{\oplus 3})_{1}$ $(AD_{3\frac{1}{2}})_{1}$ $(D_{4})_{1}$ 1
1. We determine the 2d RCFT associated with each entry of the extended magic triangle at level one. The characters of these RCFTs satisfy a uniform two-parameter family of fourth-order holomorphic modular linear differential equations (MLDEs), given in (41).
@@
2. The magic coset (48) provides a natural generalization of the dual pair structure with respect to $(E_{8})_{1}$ in the Cvitanović–Deligne exceptional series to the entire magic triangle.
@@
3. The dimensions, degeneracies, and modular $S$-matrices of all theories in the magic triangle are given uniformly.
@@
4. We show that all thirty theories on the level-one magic triangle can be decomposed into five elementary building blocks ${\rm Vir}^{\rm eff}_{5,2}$, ${\rm Vir}^{\rm eff}_{5,3}$, $(U_{1})_{3}$, ${\rm Vir}^{\rm e}_{6,5}$ and $D_{\rm 2A}$.
@@
5. The subexceptional series at level two exhibits $\mathcal{N}=1$ supersymmetry. The corresponding Neveu–Schwarz sector characters satisfy a one-parameter family of fourth-order fermionic MLDEs.
@@
6. Uniform coset constructions for $(\mathfrak{g}_{\rm CD})_{2}/{(\mathfrak{g}_{\rm sub})_{2}\times(A_{1})_{2}}$ and $(\mathfrak{g}_{\rm sub})_{2}/{(\mathfrak{g}_{\rm Severi})_{2}\times(U_{1})_{6}}$.
@@
This paper is organized as follows. In Section 2, we review background material on exceptional series, the magic square, and the magic triangle, and provide a brief introduction to characters of 2d RCFTs and modular linear differential equations (MLDEs). In particular, we recall the seminal result of Mathur, Mukhi, and Sen Mathur et al. (1988) that the CFT-type solutions of second-order holomorphic MLDEs are precisely the level-one affine characters of the Cvitanović–Deligne exceptional series. In Section 3, we extend the framework of Mathur–Mukhi–Sen to the full magic triangle and investigate several structural properties of this generalization, which may be regarded as an affine realization of the magic triangle at level one. Section 4 is devoted to the magic triangle at level two, with particular emphasis on the exceptional series appearing in the lower rows which exhibits more symmetries. In Section 5, we discuss the magic triangle at arbitrary higher levels. In the end, we give some concluding remarks. In Appendix A, we explicitly present the data for the extended magic triangle and CFTs at level one. In Appendix B, we explicitly give the S-matrices for the CFTs at level one of CD exceptional series. In Appendix C, we discuss the spurious solutions of the 4th order MLDE for CD exceptional series.
@@
We use $()^{\rm e}$ for a non-diagonal modular invariant of RCFT, i.e. a VOA extension. For example, ${\rm Vir}^{\rm e}_{6,5}$ is a non-diagonal modular invariant of Virasoro minimal model ${\rm Vir}_{6,5}$ that describes the critical three-state Potts model. We denote the RCFT associated to the triangle element $(\mu,\nu)$ at level $k$ as $\mathcal{T}^{(\mu,\nu)}_{k}$. Especially we denote the level-one theory as $\mathcal{T}^{(\mu,\nu)}$ for simplicity. We use subscript to denote the degeneracy of primary fields. We use $\rm eff$ to denote the effective theory of non-unitary minimal models.
2 Review on the magic triangle
@@ 2.1 Exceptional series and magic triangle
They correspond to the following extended Cvitanović–Deligne exceptional series respectively
@@
Here $e$ is a trivial element. For the simple Lie algebras $\mathfrak{g}$ in the series, which constitute the original Cvitanović–Deligne exceptional series, the dimension and the dual Coxeter number Cvitanovic (1981, 2008); Deligne (1996) are given by $\nu$ as
@@
The original CD exceptional series is extended to include intermediate Lie algebras $A_{1/2}$ and $E_{7+1/2}$, whose dual Coxeter number and dimension obey the same formula. To be specific, $A_{1/2}$ has dual Coxeter number $3/2$ and dimension $1$. In many senses, one can regard $A_{1/2}$ as Lie superalgebra $osp(1|2)$ which has dual Coxeter number $3/2$ and superdimension $1$. Then affine characters of $(A_{1/2})_{k}$ are just the supercharacters of $osp(1|2)_{k}$. An alternative interpretation is to regard $A_{1/2}$ as a formal object, denoted $LY$, such that the associated theory $(LY)k$ realizes the Lee–Yang model at level $k$, namely the effective Virasoro minimal model $\mathrm{Vir}^{\rm eff}_{(2k+3,2)}$ Duan et al. (2022). This correspondence follows from the fact that the supercharacters of $\mathfrak{osp}(1|2)_{k}$ agree, as $q$-series, with the characters of $\mathrm{Vir}^{\rm eff}_{(2k+3,2)}$. The $E_{7+1/2}$ is a famous intermediate Lie algebra with dual Coxeter number $24$ and dimension $190$. Its existence was noted both in physics Mathur et al. (1988); Cvitanovic (2008) and mathematics Cohen and de Man (1996), and it was subsequently constructed by Landsberg and Manivel using sextonions Landsberg and Manivel (2006).
@@
The magic triangle is a two-dimensional symmetric generalization of the Cvitanović–Deligne exceptional series, parameterized by a pair $(\mu,\nu)$ with $\mu\nu\geq 1$. In the case $\mu\nu=1$, all entries reduce to trivial elements. One systematic construction of the magic triangle proceeds by placing the Cvitanovi’c–Deligne exceptional series along the bottom row and the rightmost column, and then defining each remaining entry as the centralizer of the corresponding row element inside the column element Deligne and Gross (2002). The original magic triangle introduced in (Cvitanovic, 1981, Table 6) is reproduced in Table 1. The subarray with $\mu,\nu>1$, highlighted in Table 1, is known as the Freudenthal–Tits magic square. The Lie algebras appearing in the magic square admit a uniform construction in terms of pairs of normed division algebras, $\mathfrak{m}(\mathbb{A}_{1},\mathbb{A}_{2})$ for $\mathbb{A}=\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$. This construction realizes the idea that all five exceptional Lie algebras are related to the octonions; see Baez (2002) for a comprehensive review. Upon including the sextonion algebra $\mathbb{S}$, the magic square can be further extended to incorporate several intermediate Lie algebras; we refer to Lee et al. (2024b) for a detailed discussion. In this context, it is sometimes convenient to replace the parameter $\nu$ by $a=2(\nu-1)$, which takes values in the set $\{-8/5,-3/2,-4/3,-1,-2/3,0,1,2,4,6,8\}.$ Here the numbers $1,2,4,6,8$ are the real dimension of $\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{S},\mathbb{O}$. Finally, it is worth noting that the magic square has also appeared in the study of supergravity theories, see for example Anastasiou et al. (2014).
@@
For the element ${\mathfrak{g}}^{(\mu,\nu)}$ in the original magic triangle in Table 1, the dimension and dual Coxeter number have the following universal formulas
@@
The dimension formula coincides with that of the magic square derived in Landsberg and Manivel (2006). In particular, by specializing to $\mu=5$, one recovers the expressions for the dimensions and dual Coxeter numbers of the Cvitanović–Deligne exceptional series. The original magic triangle of Cvitanović and of Deligne–Gross does not include the cases $\mu,\nu=1/4$ or $4$. Incorporating these values leads naturally to an extended magic triangle. Remarkably, the universal formulas (5) and (6) continue to hold throughout this extension. For $\nu=1/4$, the extension introduces a single additional element, namely $A_{1/2}$. For $\nu=4$, one obtains the intermediate exceptional series, whose dimensions and dual Coxeter numbers were determined in Lee et al. (2024b) and are consistent with the general formulas (5) and (6). The dual Coxeter numbers and dimensions of all entries in the extended magic triangle are summarized in Table 5.
Several entries of the extended magic triangle require additional clarification. The formal groups ${\mathfrak{g}}^{(3/2,1)}$ and ${\mathfrak{g}}^{(2,2/3)}$ both have vanishing dimension and dual Coxeter number equal to 1-1. In Deligne and Gross (2002), these entries are realized concretely by the finite groups $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$, respectively. The formal group ${\mathfrak{g}}^{(4,1/3)}$ has dimension $1$ and dual Coxeter number $2/3$. It may be interpreted as a formal object, denoted $IM$, such that the associated theory $(IM)_{k}$ realizes the effective minimal model $\mathrm{Vir}^{\rm eff}_{(3k+2,3)}$ Lee et al. (2024b). There also exist three entries whose underlying Lie algebras have standard dimensions but nonstandard dual Coxeter numbers. These are ${\mathfrak{g}}^{(3,2/3)}=A_{1}$, ${\mathfrak{g}}^{(3/2,3/2)}=A_{1}$ and ${\mathfrak{g}}^{(2,3/2)}=A_{2}$. According to the general formula (6), their dual Coxeter numbers are given by $2/3$, $1/4$, and $3/2$, respectively. Equivalently, the invariant bilinear forms on these Lie algebras are rescaled by factors of $3$, 88, and $2$. Consequently, when constructing the associated Wess–Zumino–Witten models, the levels must be rescaled by the same factors, namely $3$, 88, and $2$, respectively.

### v2

@@ Abstract
The magic triangle due to Cvitanović and Deligne–Gross is an extension of the Freudenthal–Tits magic square of semisimple Lie algebras. In this paper, we identify all two-dimensional rational conformal field theories associated to the magic triangle. These include various Wess–Zumino–Witten (WZW) models, Virasoro minimal models, compact bosons and their non-diagonal modular invariants. At level one, we uncover a two-parameter family of fourth-order modular linear differential equation whose solutions yield the affine characters of all elements in the magic triangle. We further establish a universal coset relation for the whole triangle, generalizing the dual-pair structure with respect to $(E_{8})_{1}$ in the Cvitanović–Deligne exceptional series. This coset structure determines the dimensions and degeneracies of all primary fields and leads to five atomic models from which all theories in the triangle can be constructed. At level two, we find that a distinghuished row of the triangle – the subexceptional series– exhibits emergent $N=1$ supersymmetry. The corresponding Neveu–Schwarz/Ramond characters satisfy a one-parameter family of fermionic modular linear differential equations. In addition, we find several new uniform coset constructions involving WZW models at higher levels.
@@ 1 Introduction
The magic triangle of groups was discovered in both theoretical physics Cvitanovic (1981) and mathematics Deligne and Gross (2002). It may be viewed as a triangular extension of the Freudenthal–Tits magic square $\mathfrak{m}(\mathbb{A}_{1},\mathbb{A}_{2})$ of semisimple Lie algebras and Lie groups, where $\mathbb{A}$ ranges over the four normed division algebras: the real numbers, complex numbers, quaternions, and octonions. From another perspective, the magic triangle can be regarded as a two-dimensional generalization of the celebrated Cvitanović–Deligne exceptional series Cvitanovic (1981, 2008); Deligne (1996). The original magic triangle of Lie groups, found by Cvitanović (Cvitanovic, 1981, Table 6) from certain Diophantine conditions on representation dimensions in the context of extended supergravity theories, is reproduced in Table 1. The magic triangle of Deligne and Gross is built as a system of centralizers from dual pairs in $E_{8}$, see (Deligne and Gross, 2002, Table 1).
The magic triangle is symmetric, and each row (or, equivalently, each column) is referred to as an exceptional series. In Table 1, the bottom row, terminating at $E_{8}$, is known as the Cvitanović–Deligne exceptional series. The row immediately above it, ending at $E_{7}$, is called the subexceptional series. The third row from the bottom, ending at $E_{6}$, is known as the Severi exceptional series, while the fourth row from the bottom, terminating at $F_{4}$, is referred to as the Severi-section exceptional series.
@@
Each exceptional series exhibits distinctive and remarkable properties in representation theory. For instance, the Cvitanović–Deligne exceptional series displays striking uniformity in the tensor product decompositions of adjoint representations, and the dimensions of its irreducible representations admit expressions as rational functions of the dual Coxeter number Deligne (1996); Cohen and de Man (1996); Deligne and de Man (1996). This series appears frequently in theoretical physics, including 2d rational CFTs Mathur et al. (1988), 4d $N=2$ superconformal field theories (SCFTs) (see e.g. (Beem and Rastelli, 2018, Section 4)), and 6d $N=(1,0)$ SCFTs Morrison and Taylor (2012). The subexceptional series first appeared as global symmetries in 4d extended supergravity theories Cvitanovic (1981), and it also possesses many distinguished representations with uniform tensor product decompositions and dimension formulas Landsberg and Manivel (2004). From the perspective of Vogel’s theory of the universal Lie algebra, both the Cvitanović–Deligne series and the subexceptional series form straight lines in Vogel’s projective plane Vogel (1999). Moreover, the four exceptional series are deeply connected with geometry Landsberg and Manivel (2002, 2004, 2001). In particular, the terminology Severi exceptional series and Severi-section exceptional series reflects their geometric relationship with Severi varieties Landsberg and Manivel (2004).
@@
Several conceptual issues arise in the original magic triangle displayed in Table 1. First, despite its name, the original magic triangle of Lie groups does not form a complete triangular array, as certain entries along the slope are missing. These absent elements were identified as discrete groups in Deligne and Gross (2002). A second issue concerns the common practice in the physics literature of adjoining an additional entry “Lee–Yang model” at the beginning of the Cvitanović–Deligne exceptional series Mathur et al. (1988); Beem and Rastelli (2018), which further obscures the triangular structure. A third question is whether the famous intermediate Lie algebra $E_{7+1/2}$ Landsberg and Manivel (2006), lying between $E_{7}$ and $E_{8}$, can be incorporated into the Cvitanović–Deligne exceptional series in a manner compatible with the magic triangle. In this work, we address these three issues in a unified framework and show that their resolution naturally requires an extension of the original diagram, resulting in a genuinely triangular structure.
@@
Our approach is based on two-dimensional rational conformal field theories (RCFTs) associated with the magic triangle. This may be viewed as a natural generalization of the work of Mathur, Mukhi, and Sen on the 2d RCFTs associated with the Cvitanović–Deligne exceptional series Mathur et al. (1988, 1989). From this perspective, the resulting structure can be interpreted as an affine analogue of the original magic triangle. The construction is particularly natural at level one; the resulting extended triangle is displayed in Table 2. For each Lie group appearing in the original magic triangle in Table 1, the corresponding entry in Table 2 is given by the associated Wess–Zumino–Witten model mostly at level one. The remaining entries consist of various minimal models, compact boson theories, and their non-diagonal modular invariants, which will be explained in detail below. In particular, we introduce an additional row ending at $E_{7+1/2}$, referred to as the intermediate exceptional series, which has been studied systematically in Lee et al. (2024b).
@@ 0 $D_{2\rm A}$ $M_{2}$ $(A_{1}^{\oplus 3})_{1}$ $(AD_{3\frac{1}{2}})_{1}$ $(D_{4})_{1}$ 1
1. We determine the 2d RCFT associated with each entry of the extended magic triangle at level one, as presented in Table 2. The characters of these RCFTs satisfy a uniform two-parameter family of fourth-order holomorphic modular linear differential equations (MLDEs), given in (41).
@@
2. The magic coset (48) provides a natural generalization of the dual-pair structure with respect to $(E_{8})_{1}$ in the Cvitanović–Deligne exceptional series to the entire magic triangle.
@@
3. The dimensions, degeneracies, and modular $S$-matrices of all theories in the magic triangle at level one are given in a uniform manner.
@@
4. We show that all thirty theories on the level-one magic triangle can be decomposed into five elementary building blocks: Vir^eff_5,2, Vir^eff_5,3, (U_1)_3, Vir^e_6,5, D_2A.
@@
5. The subexceptional series at level two exhibits N=1{N}=1 supersymmetry. The corresponding fermionic characters satisfy a one-parameter family of fourth-order fermionic MLDEs.
@@
6. We identify many uniform coset constructions involving exceptional series at level two such as $(\mathfrak{g}_{\rm CD})_{2}/{(\mathfrak{g}_{\rm sub})_{2}\times(A_{1})_{2}}$ and $(\mathfrak{g}_{\rm sub})_{2}/{(\mathfrak{g}_{\rm Severi})_{2}\times(U_{1})_{6}}$.
@@
This paper is organized as follows. In Section 2, we review background material on exceptional series, the magic square, and the magic triangle, and provide a brief introduction to characters of 2d RCFTs and modular linear differential equations (MLDEs). In particular, we recall the seminal result of Mathur, Mukhi, and Sen Mathur et al. (1988) that the CFT-type solutions of second-order holomorphic MLDEs are precisely the level-one affine characters of the Cvitanović–Deligne exceptional series. In Section 3, we extend the framework of Mathur–Mukhi–Sen to the full magic triangle and investigate several structural properties of this generalization, which may be regarded as an affine realization of the magic triangle at level one. Section 4 is devoted to the magic triangle at level two, with particular emphasis on the exceptional series appearing in the lower rows which exhibits richer symmetry structures. In Section 5, we discuss the magic triangle at arbitrary positive integer levels. In the end, we give some concluding remarks. In Appendix A, we explicitly present the data for the extended magic triangle and the CFTs at level one. In Appendix B, we explicitly give the $S$-matrices for the CFTs at level one of Cvitanović–Deligne exceptional series. In Appendix C, we discuss the spurious solutions of the 4th order MLDE for Cvitanović–Deligne exceptional series.
@@
We use $()^{\rm e}$ for a non-diagonal modular invariant of RCFT, that is, a VOA extension. For example, ${\rm Vir}^{\rm e}_{6,5}$ is a non-diagonal modular invariant of Virasoro minimal model ${\rm Vir}_{6,5}$ that describes the critical three-state Potts model. We denote the RCFT associated with the triangle element $(\mu,\nu)$ at level $k$ as $\mathcal{T}^{(\mu,\nu)}_{k}$. In particular, we denote the level-one theory as $\mathcal{T}^{(\mu,\nu)}$ for simplicity. We use subscript numbers to denote the degeneracies of primary fields, and we use the superscript $\rm eff$” to denote the effective theory of non-unitary minimal models.
2 Review of the magic triangle
@@ 2.1 Exceptional series and magic triangle
These values correspond to the following extended Cvitanović–Deligne exceptional series, respectively:
@@
Here $e$ denotes a trivial element. For the simple Lie algebras $\mathfrak{g}$ in the series, which constitute the original Cvitanović–Deligne exceptional series, the dimension and the dual Coxeter number Cvitanovic (1981, 2008); Deligne (1996) are expressed in terms of $\nu$ as
@@
The original Cvitanović–Deligne exceptional series can be extended to include intermediate Lie algebras $A_{1/2}$ Shtepin (1993) and $E_{7+1/2}$ Landsberg and Manivel (2006), whose dual Coxeter number and dimension obey the same formulas (3) and (4). [^1] The notation “half” indicates that the dual Coxeter number of the intermediate Lie algebra is the arithmetic mean of those of the two algebras appearing in the corresponding inclusions, namely $A_{0}\subset A_{1/2}\subset A_{1}$ and $E_{7}\subset E_{7+1/2}\subset E_{8}$. To be specific, $A_{1/2}$ has dual Coxeter number $3/2$ and dimension $1$. In many senses, one may regard $A_{1/2}$ as Lie superalgebra $osp(1|2)$ which has dual Coxeter number $3/2$ and superdimension $1$. Then affine characters of $(A_{1/2})_{k}$ are just the supercharacters of $osp(1|2)_{k}$. An alternative interpretation is to regard $A_{1/2}$ as a formal object, denoted $LY$, such that the associated theory $(LY)_{k}$ realizes the Lee–Yang model at level $k$, namely the effective Virasoro minimal model $\mathrm{Vir}^{\rm eff}_{(2k+3,2)}$ Duan et al. (2022). This correspondence follows from the fact that the supercharacters of $\mathfrak{osp}(1|2)_{k}$ agree, as $q$-series, with the characters of $\mathrm{Vir}^{\rm eff}_{(2k+3,2)}$ Ferrari et al. (2024). On the other hand, the $E_{7+1/2}$ is a famous intermediate Lie algebra with dual Coxeter number $24$ and dimension $190$. Its existence was noted both in physics Mathur et al. (1988); Cvitanovic (2008) and mathematics Cohen and de Man (1996), and it was subsequently constructed by Landsberg and Manivel using sextonions Landsberg and Manivel (2006). Many irreducible representations of $E_{7+1/2}$ have been studied in Lee et al. (2024a). The affine characters of $(E_{7+1/2})_{1}$ were found both by MLDE Mathur et al. (1988) and Hecke operator Harvey and Wu (2018). The affine characters of $(E_{7+1/2})_{2}$ were recently found by Hecke operator in Duan et al. (2022).
@@
The magic triangle is a two-dimensional symmetric generalization of the Cvitanović–Deligne exceptional series, parameterized by a pair $(\mu,\nu)$ with $\mu\nu\geq 1$. In the case $\mu\nu=1$, all entries reduce to trivial elements due to the zero dimension. [^2] They may nevertheless be interesting, and we hope to return to this point elsewhere. One systematic construction of the magic triangle proceeds by placing the Cvitanović–Deligne exceptional series along the bottom row and the rightmost column, and then defining each remaining entry as the centralizer of the corresponding row element inside the column element Deligne and Gross (2002). The original magic triangle introduced in (Cvitanovic, 1981, Table 6) is reproduced in Table 1. The subarray with $\mu,\nu>1$, highlighted in Table 1, is known as the Freudenthal–Tits magic square. The Lie algebras appearing in the magic square admit a uniform construction in terms of pairs of normed division algebras, denoted by $\mathfrak{m}(\mathbb{A}_{1},\mathbb{A}_{2})$ for $\mathbb{A}=\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$. This construction realizes the idea that all five exceptional Lie algebras are related to the octonions; see Baez (2002) for a comprehensive review. Upon including the sextonion algebra $\mathbb{S}$, the magic square can be further extended to incorporate several intermediate Lie algebras; see Westbury (2006); Lee et al. (2024b) for a detailed discussion. In this context, it is sometimes convenient to replace the parameter $\nu$ by $a=2(\nu-1)$, which takes values in the set $\{-8/5,-3/2,-4/3,-1,-2/3,0,1,2,4,6,8\}.$ Here the numbers $1,2,4,6$, and $8$ are the real dimensions of $\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{S}$, and $\mathbb{O}$, respectively. Finally, it is worth noting that the magic square has also appeared in the study of supergravity theories, see for example Anastasiou et al. (2014).
@@
For the element ${\mathfrak{g}}^{(\mu,\nu)}$ in the original magic triangle in Table 1, the dimension and dual Coxeter number admit the following universal formulas:
@@
The dimension formula was given in (Cvitanovic, 1981, Equation (11.1)) in a different form, and also given for magic square in (Landsberg and Manivel, 2006, Proposition 3.2). In particular, by specializing to $\mu=5$, one recovers the expressions for the dimensions (3) and dual Coxeter numbers (4) of the Cvitanović–Deligne exceptional series. The original magic triangle of Cvitanović and Deligne–Gross does not include the cases $\mu,\nu=1/4$ or $4$. Incorporating these values naturally leads to an extended magic triangle. Interestingly, the universal formulas (5) and (6) continue to hold throughout this extension. In particular, the dimensions are all non-negative integers. For $\nu=1/4$, the extension introduces a single additional element, namely $A_{1/2}$. For $\nu=4$, one obtains the intermediate exceptional series, whose dimensions and dual Coxeter numbers were determined in Lee et al. (2024b) and are consistent with the general formulas (5) and (6). The dual Coxeter numbers and dimensions of all entries in the extended magic triangle are summarized in Table 5.
Several entries of the extended magic triangle require additional clarification. The formal groups ${\mathfrak{g}}^{(3/2,1)}$ and ${\mathfrak{g}}^{(2,2/3)}$ both have vanishing dimension and dual Coxeter number equal to 1-1. In Deligne and Gross (2002), these entries are realized concretely by the finite groups $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$, respectively. The element ${\mathfrak{g}}^{(4,1/3)}$ has dimension $1$ and dual Coxeter number $2/3$. It may be interpreted as a formal object, denoted $IM$, such that the associated theory $(IM)_{k}$ realizes the effective minimal model $\mathrm{Vir}^{\rm eff}_{(3k+2,3)}$ Lee et al. (2024b). There also exist three entries whose underlying Lie algebras have standard dimensions but nonstandard dual Coxeter numbers. These are ${\mathfrak{g}}^{(3,2/3)}=A_{1}$, ${\mathfrak{g}}^{(3/2,3/2)}=A_{1}$ and ${\mathfrak{g}}^{(2,3/2)}=A_{2}$. According to the general formula (6), their dual Coxeter numbers are given by $2/3$, $1/4$, and $3/2$, respectively. This means that the invariant bilinear forms on these Lie algebras are rescaled by factors of $3$, 88, and $2$. Consequently, when constructing the associated Wess–Zumino–Witten models, the levels must be rescaled by the same factors, namely $3$, 88, and $2$, respectively.